five less than a number is 15. Find the number.
Answer by actuary(112)

Let n = the number you are interested in finding. The description of the problem says "5 less than the number is 15". The key is to find the important words such as "is less than" and "is". "Is less than" means "subtraction" and "is" mean "= (equals)" Rewriting the words as an algebraic equation you have n-5=15. Using the "Rules of algebra", add 5 to each side of the equation, s0 n-5+5=15+5. So n=20.
Good luck studying math.
Larry